A gas absorbs 3.4 kJ of heat and does 1.9 kJ of work. Calculate ΔE.
1
Expert's answer
2020-02-26T05:57:26-0500

First law of thermodynamics:

"\Delta E= Q-W
\\ \Delta E= 3.4-1.9=1.5 \ kJ"